A mechanic can install carburetors in 3 cars every 4 hours. At that rate, how long will it take the mechanic to install carburetors in 5 cars?

Answer by checkley77(12844)

3/4=5/X
3X=20
X=20/3
X=6.667 HOURS ANS.